OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y - z = 2 3x + y - 2z = -15 3x + 4y - 3z = -2 Choose one answer. A. (7, -4, -1) B. (-7, -4, 1) C. (7, -4, 1) D. (-7, 4, -1)

OpenStudy (anonymous):

start substituting son

OpenStudy (anonymous):

Help? I'm giving medals

OpenStudy (anonymous):

oh gosh medals.

OpenStudy (anonymous):

start it first, ill tell you where you go wrong

OpenStudy (texaschic101):

x + 2y - z = 2 x = -2y + z + 2 3x + y - 2z = -15 3(-2y + z + 2) + y - 2z = -15 -6y + 3z + 6 + y - 2z = -15 -5y + z = -15 - 6 -5y + z = - 21 z = 5y - 21 3x + 4y - 3z = -2 3(-2y + z + 2) + 4y - 3z = -2 -6y + 3z + 6 + 4y - 3z = -2 -2y = -2 - 6 -2y = -8 y = 4 z = 5y - 21 z = 5(4) - 21 z = 20 - 21 z = -1 x = -2y + z + 2 x = -2(4) + (-1) + 2 x = -8 -1 + 2 x = -9 + 2 x = -7